OpenStudy (el_arrow):

help me find the derivative of x3^x/3x-1

OpenStudy (abdullah1995):

x3^x ?

OpenStudy (el_arrow):

do i bring the x down or what?

OpenStudy (el_arrow):

yeah

OpenStudy (el_arrow):

sorry i mean (x3^x)/3^(x)-1

OpenStudy (anonymous):

\[\frac{d}{dx}\left[\frac{x3^x}{3^x-1}\right]\]

OpenStudy (abdullah1995):

is that right ^

OpenStudy (el_arrow):

yep that correct

OpenStudy (abdullah1995):

have you studied ln ?

OpenStudy (el_arrow):

yeah but i still dont get it :(

OpenStudy (anonymous):

Some tools you'll need: \[\textbf{Product rule}\\\textbf{Quotient rule}\\ \textbf{Derivatives of exponential function}\text{, i.e. }\frac{d}{dx}e^x=e^x\]

OpenStudy (el_arrow):

so you use the quotient rule and e on this one?

OpenStudy (abdullah1995):

e^x = e^x is special case

OpenStudy (abdullah1995):

ok it's really simple

OpenStudy (anonymous):

I mean for bases other than \(e\), you can still use the derivative form of \(e^x\). \[b^x=e^{\ln b^x}=e^{x\ln b}~~\implies~~\frac{d}{dx}b^x=\ln b\cdot e^{x\ln b}=\ln b\cdot b^x\]

OpenStudy (el_arrow):

so it would be 3e^x

OpenStudy (abdullah1995):

https://awwapp.com/draw.html#ce7077cd

OpenStudy (abdullah1995):

click on the link it will lead you to whiteboard. we will solve it there

OpenStudy (abdullah1995):

let me know when you are in

OpenStudy (el_arrow):

i am there